QUESTION 3 How many millimoles of NaOH will react completely with 25 mL of 1.5 M H2C2047 H2C204 + 2 NaOH = Na2C204 + 2H20
Answer #1

ANSWER- milimoles of NaOH= 75

Solution-

Equation for reaction b/w NaOH and oxalic acid(C2H2O4) is-

2 NaOH + C2H2O4 = C2O4Na2 + 2 H2O

Molar mass of NaOH = 23 + 16 + 1 = 40 g/mole

Molar mass of C2H2O4 = 24 + 2 + 64 = 90 g/mole

As 2 mole of NaOH reacts with 1 mole oxalic acid,

Milimoles of C2H2O4 = volume (in mL) × Molarity

Milimoles of C2H2O4 = 25 x 1.5 = 37.5

Hence, milimoles of NaOH = 2 × milimoles of C2H2O4 = 2 × 37.5 = 75 milimoles